OpenStudy (anonymous):

When rolling two number cubes or a set of six-sided dice, what is the probability of rolling a 7? A. 1/36 B. 1/18 C. 1/6 D. 1/3

OpenStudy (anonymous):

A

OpenStudy (anonymous):

????

OpenStudy (raden):

hint : events of 7 = (1,6),(6,1),(2,5),(5,2),(3,4),(4,3). there are 6 events and we knowed that all event = 6^2 = 36 so, what's your answer ?

OpenStudy (anonymous):

A

OpenStudy (raden):

probability = event/all event it is 6/36 = 1/6

OpenStudy (anonymous):

OKAY I WILL TRY THAT ONE THANKS AND I GIVE U A MEDAL :)

OpenStudy (raden):

you're welcome :)